### v1

# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

Abstract.
As algorithmic systems increasingly structure platform labor, workers often rely on informal “folk theories”, experience-based beliefs about how algorithms work, to navigate opaque and unstable algorithmic environments. Prior research has largely treated these theories as bottom-up, peer-driven strategies for coping with algorithmic opacity and uncertainty. In this study, we shift analytical attention to intermediary organizations and examine how folk theories of algorithms can be institutionally constructed and operationalized by those organizations as tools of labor management. Drawing on nine months of ethnographic fieldwork and 37 interviews with live-streamers and staff at Multi-Channel Networks (MCNs) in China, we show that MCNs develop and circulate dual algorithmic theories: internally, they acknowledge the volatility of platform systems and adopt probabilistic strategies to manage risk; externally, they promote simplified, prescriptive theories portraying the algorithm as transparent, fair, and responsive to individual effort. They have further operationalize those folk theories for labor management, encouraging streamers to self-discipline and invest in equipment, training, and routines, while absolving MCNs of accountability. We contribute to CSCW and platform labor literature by demonstrating how informal algorithmic knowledge, once institutionalized, can become infrastructures of soft control – shaping not only how workers interpret platform algorithms, but also how their labor is structured, moralized and governed.
Algorithm; Platform; Folk Theories of Algorithms; Live-streaming; MCNs; China; Ethnography
This work was completed during Rongyi Chen’s visiting research at George Mason University.
This work was completed during Xianzhe Fan’s visiting research at Carnegie Mellon University.
$c$s: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
As algorithmic systems become increasingly central to digital labor, workers across domains, ride-hailing, content moderation, e-commerce, and live-streaming, are compelled to navigate opaque, volatile, and often inscrutable mechanisms of evaluation and visibility. Algorithms do not simply mediate tasks; they embed managerial logics into platform infrastructure, determining what is seen, who is rewarded, and how performance is measured (Wood et al., 2019; Bucher, 2018; Kellogg et al., 2020). For workers, this opacity creates a pervasive sense of precarity: their success of the work appears contingent not only on effort or skill, but on deciphering hidden systems that govern algorithmic recognition. In response, scholars in HCI, CSCW, and media studies have increasingly turned their attention to the interpretive practices workers develop to make sense of algorithmic environments. A key concept emerging from this line of research is folk theories of algorithms—informal beliefs based on personal experience about how algorithms function (DeVito et al., 2017; Xiao et al., 2024b; Ytre-Arne and Moe, 2021). These theories allow workers to construct actionable understandings of platform behavior, guiding decisions about when, how, and what to produce. Prior work has richly documented how workers share folk theories in peer networks (McDaid et al., 2023; Yao et al., 2021), use them to rationalize inconsistent outcomes (Christin, 2020; Poon et al., 2021), or deploy them to strategically align with perceived algorithmic preferences (Bonini and Treré, 2024; Lei, 2021). More recent scholarship also examines how these theories are not developed in isolation but are shaped collectively within communities (Xiao et al., 2024b; Zhang et al., 2023). However, this growing body of work continues to frame folk theories primarily as bottom-up epistemic practices—emergent, community-driven, and reactive. What remains underexplored is how such theories can be deliberately organized, curated, and operationalized by third-party institutions. That is, how folk theories shift from informal peer-to-peer artifacts to formalized narratives embedded in managerial routines, spatial arrangements, and incentive structures.
This paper addresses this gap by examining the role of Multi-Channel Networks (MCNs) in China’s live-streaming industry. MCNs are intermediary organizations that operate between individual streamers and platform companies. They help streamers get started, improve their content, build their personal brand, and make money, but also often control what streamers produce and how they work (Zhang et al., 2024; Zhang and Tong, 2024; Liang and Ji, 2024). While MCNs have been widely studied in western contexts such as YouTube as service providers that support monetization, branding, and creator management (Siciliano, 2023; Gardner and Lehnert, 2016), these research often overlook their epistemic role, especially in algorithm interpretation. We advance this literature by theorizing MCNs not just as economic intermediaries, but as organizational actors that actively interpret and institutionalize algorithmic knowledge. This shift foregrounds how algorithmic sensemaking is produced and circulated at scale. China’s live-streaming industry then offers a particularly generative setting for this analysis. Unlike more entertainment-driven and part-time creator cultures elsewhere, Chinese live-streaming is highly industrialized and labor-intensive (Ye et al., 2023; Sun et al., 2023), with MCNs exerting deep control over labor and visibility. Despite their waning influence on Western markets (Alexander, 2018), MCNs remain dominant in China, shaping how streamers understand and respond to algorithmic systems. Given the generally low algorithmic literacy of streamers (Xiao et al., 2024a; Tang et al., 2022), MCNs possess disproportionate epistemic authority, determining not just what is knowable about algorithms, but who gets to know it and how.
In this study, we focus on how Chinese MCNs develop and deploy folk theories of algorithms organizationally for labor management in live-streaming industry. Drawing on nine months of ethnographic fieldwork and 37 interviews with MCN staff and streamers, we ask:
  • RQ1: How do MCNs interpret and shape streamers’ understanding of algorithms on Chinese live-streaming platforms?
  • RQ2: How are these interpretations translated into everyday organizational practices for labor management?
Rather than treating folk theories of algorithms as bottom-up sensemaking tools, our findings reveal how they are deliberately constructed, circulated, and embedded by intermediary organizations (MCNs) to govern labor under algorithmic opacity. Specifically, we find that MCNs construct dual folk theories of algorithms: internally, they acknowledge the instability and opacity of platform algorithms, using this uncertainty to manage organizational risk; externally, they promote motivational narratives that frame the algorithm as fair, responsive, and effort-driven, shifting the burden of performance onto individual streamers. These folk theories are institutionalized through moral discourse, performance rankings, aesthetic standards, spatial control, and financial dependencies, making speculative knowledge of algorithms feel actionable and enforceable.
This study offers three key contributions to the CSCW/HCI literature. First, it re-theorizes folk algorithmic knowledge as institutionalized epistemic infrastructure, not merely emergent and experiential, but strategically produced and maintained within organizations. Second, it conceptualizes these folk theories as tools of algorithmic labor management, enabling MCNs to govern workers through belief, aspiration, and moral obligation. Third, it provides design and policy implications for CSCW and HCI, urging system designers and policy makers to account for epistemic authority that mediates algorithmic work.
2. Background and Related Work
This section reviews three strands of literature that form the foundation of our study. First, we examine research on folk theories of algorithms (subsection 2.1). Second, we analyze the shift from direct algorithmic management (e.g., task allocation or performance metrics) to softer and more affective forms of control (subsection 2.2). Third, we turn to literature on Multi-Channel Networks (MCNs) and surrounding studies on the live-streaming industry (subsection 2.3).
2.1. Folk Theories of Algorithms
In this section, we examine two interrelated dimensions around algorithms. First, we explore how workers, both individually and collectively, construct informal algorithmic knowledge to adapt and survive amid opacity and volatility (subsubsection 2.1.1). Second, we investigate how organizations actively shape and institutionalize folk theories to align worker behavior with algorithmic logics and organizational goals (subsubsection 2.1.2).
2.1.1. From Individual to Collective Folk Theories of Algorithms
As algorithmic systems have grown increasingly opaque, unstable, and consequential, users and workers have developed informal interpretive frameworks, often termed folk theories of algorithms, to perceive, understand and make sense of how these systems function (DeVito et al., 2017; Bucher, 2018; Dogruel, 2021; Huang et al., 2022). Rooted in everyday experiences rather than technical knowledge, such folk theories reflect individuals’ efforts to explain, predict, and act upon algorithmic outcomes in the absence of transparent information, often drawing on trial and error, peer advice, or rumor  (Ytre-Arne and Moe, 2021; Siles et al., 2020), while most studies frame folk theories as enabling users to preserve a sense of comprehension and control despite the opacity of algorithmic systems  (Eslami et al., 2016; Rader and Gray, 2015).
In the context of platform labor, folk theories also serve as practical guides among workers to navigating algorithmic systems, such as those that determine content visibility or task allocation (McDaid et al., 2023; Bishop, 2020, 2019; DeVito, 2022). Workers on platforms like YouTube  (Bishop, 2020, 2019), TikTok  (DeVito, 2022), or Uber  (Cameron et al., 2021b, a; Cameron, 2020) often develop personalized work patterns from folk theories that shape their daily routines, even if these are technically inaccurate. Some scholars observed that ride-share drivers, confronted with opaque algorithmic decision-making, developed personalized explanations and workaround strategies to navigate the system, which were frequently discussed and refined within online peer communities  (Tarafdar et al., 2023; Ramesh et al., 2023; Lee et al., 2015; Rosenblat, 2018). In such settings, folk theories are rarely the product of isolated reasoning. Instead, they often emerge through collective construction, a process whereby individual experiences, suspicions, and partial observations are aggregated, debated, and iteratively reinforced within worker networks  (Xiao et al., 2024b; DeVito et al., 2017; Bonini and Treré, 2024; Karizat et al., 2021). Over time, these collaboratively formed beliefs can crystallize into shared logics that inform daily labor practices, regardless of their technical accuracy. These theories are not static; they circulate dynamically across online forums, social media groups, and informal conversations, adapting to shifts in platform policies or worker experiences  (DeVito, 2021; Bonini and Treré, 2024).
Thus, folk theories of algorithms operate as collective coping mechanisms, socially constructed and socially sustained forms of knowledge that make opaque algorithmic systems actionable and survivable in practice  (Xiao et al., 2024b). Yet, their influence extends beyond individual adaptation: as they circulate, stabilize, and sediment into routine, these theories begin to shape the very structure of platform work itself. They define what counts as “smart” strategies, legitimate routines, or risky behaviors under algorithmic governance, thereby producing normative expectations that guide how workers plan, perform, and assess their labor. Over time, these collectively maintained theories can crystallize into informal occupational scripts—shared understandings about how to succeed, what to avoid, and how to interpret algorithmic feedback  (Cameron, 2024; Meijerink and Bondarouk, 2023). Such scripts inform not only individual choices but also broader cultural logics within platform economies, reinforcing ideals of hustle, responsiveness, or creative optimization  (Xiao et al., 2024b). In this sense, folk theories are not merely responses to uncertainty; they are productive forces that help structure platform labor itself, embedding behavioral norms, emotional orientations, and survival tactics into the everyday life of algorithmic work  (Kellogg et al., 2020). To fully grasp the governance potential of folk theories, we must turn to the organizational and platform forces that selectively shape, formalize, and instrumentalize these beliefs.
2.1.2. From Interpretive Folk Theories to Organizational Control
While much of the existing literature conceptualizes folk theories as grassroots knowledge emerging from lived experience, a growing body of research highlights that such theories can also be constructed, curated, and institutionalized by organizations (Siciliano, 2023; Zhang and Tong, 2024). An emerging research direction asks how platform organizations, managers, or intermediary institutions actively shape, amplify, or exploit these user beliefs as part of their strategic governance toolkit  (Cornils, 2020; Randhawa et al., 2017; Ticona, 2022; Xiao et al., 2024b).
Platforms play a foundational role in shaping user beliefs through deliberate ambiguity, selective disclosure, and interface nudges. For example, Uber, the platform’s vague deactivation policies and gamified app interface led many drivers to internalize the belief that low acceptance rates would lead to penalties, even though this was never officially confirmed (Rosenblat and Stark, 2016; Heiland, 2025). Uber did little to dispel this belief, and it arguably served the company’s interest by promoting behavioral compliance. Similarly, Ma and Kou’s study on YouTube demonstrates how opaque demonetization practices forced creators to develop their own moderation heuristics, such as avoiding controversial keywords or altering content format, to appease the perceived preferences of the monetization algorithm (Ma and Kou, 2021). In both cases, the platforms shaped the contours of folk theories by withholding definitive explanations while maintaining algorithmic authority, thereby offloading interpretive labor onto users.
As Rosenblat and Stark argue in their study of Uber, the asymmetries of information and power embedded in the platform’s technical and communicative infrastructure are central to its ability to manage workers without formal employment structures. In their view, algorithmic systems are not only functional tools but also rhetorical devices that legitimize corporate control while disempowering labor  (Rosenblat and Stark, 2016). This discursive and infrastructural opacity enables platforms to offload interpretive labor onto workers, who in turn develop folk theories that align their behavior with the platform’s operational goals, even when these theories are based on speculation. These insights form the theoretical foundation of our study: that sometimes users’ understandings of algorithms can be deliberately shaped and directed by institutional actors.
What remains underexamined, however, is how these processes unfold not within platforms themselves, but through third-party actors that mediate between platforms and workers. If platforms benefit from maintaining algorithmic ambiguity, what role do intermediary organizations play in translating, amplifying, or formalizing that ambiguity? Recent studies have begun to explore this intermediary layer. For instance, Siciliano demonstrates how YouTube-based MCNs provide training and support to creators based on speculative algorithmic interpretations, reinforcing platform logics under the guise of optimization  (Siciliano, 2023). Xiao et al. examine how fan clubs as a force beyond platforms constructed an elaborate rhetorical discourse around algorithmic operations, along with concrete manipulation tutorials, to mobilize millions of ordinary users into collective action aimed at influencing recommendation algorithms  (Xiao et al., 2024b). These works on how third-party actors beyond platforms act in algorithmic labor governance and bring attention to the sociotechnical infrastructures through which folk theories circulates. However, we still know little about how these institutions embed folk theories into material and contractual infrastructures that shape labor discipline. To address these questions, our study investigates how algorithmic uncertainty is interpreted, standardized, and embedded into institutional practices, not merely as speculation, but as an instrument of labor governance in Chinese live-streaming industry.
2.2. Algorithmic Management
Algorithmic management refers to the use of computational systems to allocate, supervise, and evaluate labor, increasingly replacing or supplementing traditional managerial functions in digital and platform-mediated work environments (Kellogg et al., 2020; McDaid et al., 2023; Möhlmann et al., 2021). Early scholarship in this field has focused on how platforms use algorithms to automate core managerial tasks, such as task assignment, performance monitoring, and discipline, especially in gig economy sectors like ride-hailing, delivery, and content moderation (Meijerink and Bondarouk, 2023; Momin, 2023). These studies highlight how algorithmic opacity and asymmetrical access to data reinforce structural inequalities, subjecting workers to rigid systems with limited transparency, accountability, or recourse (Rosenblat and Stark, 2016; Christin, 2020). Algorithms, in this view, become infrastructure for managerial control, embedding hierarchical logics in seemingly neutral technical systems.
Recent work has expanded the analytical lens beyond technical automation to include the social, affective, and discursive dynamics of algorithmic management. Scholars have documented how platforms govern not only through commands but through soft control, i.e., the subtle shaping of worker behavior via gamified dashboards, reward systems, performance metrics, and motivational narratives (Cameron, 2020; Jarrett, 2022; Wood and Lehdonvirta, 2023). These systems encourage workers to internalize platform values such as meritocracy, efficiency, and visibility, often prompting aspirational self-discipline under conditions of precarity. This turn in the literature emphasizes how algorithmic power operates at the level of subject formation, producing compliant, hopeful, and responsibilized workers even in the absence of overt coercion (Dasgupta et al., 2025; Cameron et al., 2021a).
However, much of this research has concentrated on the direct relationship between platforms and workers, with comparatively less attention to how intermediary institutions, such as third-party service providers, outsourcing firms, or managerial brokers, mediate, interpret, and operationalize algorithmic systems. These actors increasingly serve as epistemic and organizational bridges between opaque platforms and fragmented labor forces, playing a significant yet under-theorized role in shaping how algorithmic control is enacted on the ground. Our study contributes to this emerging line of inquiry by examining how such intermediaries not only interpret algorithmic logic but also strategically construct and institutionalize folk theories of algorithms to govern labor through belief, affect, and aspiration.
2.3. MCNs and Live-Streamers
This study investigates how folk theories of algorithms are strategically constructed, circulated, and institutionalized by intermediary organizations within China’s live-streaming industry. We focus on live-streaming industry as a particularly consequential site: streamers operate in highly volatile algorithmic environments, face extreme uncertainty in monetization, and often lack the technical knowledge to interpret the systems that shape their visibility and income (Sun et al., 2023; Yao et al., 2023; Hou et al., 2020; TikTok, 2022). These epistemic conditions make streamers especially dependent on externally provided explanations of how platform algorithms “work”, creating great context for the emergence and manipulation of folk theories.
Unlike traditional gig work structured around discrete tasks, live-streamers must continuously perform for audiences in real time, often late into the night, while adapting their appearance, affect, and persona to meet platform and viewer expectations (Freeman and Wohn, 2020; Momin, 2023; Ye et al., 2023). The platform environment further exacerbates this instability. Streamers must rely on opaque algorithmic recommendation systems to gain visibility, yet they have little control or understanding of how these systems function (Sun et al., 2023; Yao et al., 2023). Success is contingent on being surfaced by algorithms, but algorithmic volatility means that visibility is uneven and unpredictable. Income models are primarily gift-based: most streamers do not receive a base salary, and instead depend on voluntary tips or fan gifts, the digital transactions that function more like micro-donations than stable compensation (Hou et al., 2020; Sun et al., 2023). With platforms retaining a significant portion of these gifts, and no obligation for viewers to contribute, streamers are left to shoulder both the risks and uncertainties of monetization (TikTok, 2022). Many live-streamers, particularly in the Global South or lower-income regions, enter this labor market with limited digital literacy and educational resources (Momin, 2023; Xiao et al., 2024a; Tang et al., 2022). They often lack the technical background to interpret algorithmic behavior, analyze platform metrics, or optimize engagement strategies. This limited algorithmic understanding further amplifies their precarity, making them more vulnerable to misinformation, overwork, and self-blame when platform performance falters (Uttarapong et al., 2021; Ye et al., 2023).
In this context, MCNs have become central to the construction of algorithmic knowledge and the mediation of labor, which has been a great case for our study in institutionalized folk theories of algorithms. Originating in Western markets such as YouTube, the MCN model initially offered monetization support, branding strategy, and analytics services to individual creators and channel collectives (Lobato, 2016; Postigo, 2016). While the influence of MCNs on YouTube has fluctuated, research has shown that these entities helped consolidate and industrialize early creator labor by aggregating talent, mediating intellectual property claims, and standardizing aesthetic and technical norms across channels (Cunningham and Craig, 2019; Chen et al., 2024). Even when not formally embedded within platform governance, MCNs in the West have played significant roles in disciplining creative labor through soft control mechanisms such as branding scripts, content quotas, and algorithm-focused consultations (Siciliano, 2023).
We selected China as our main study context. In China, MCNs have become far more structurally embedded within the operational logics of platforms like Douyin, Kuaishou, and Huya. Rather than serving merely as optional support services, MCNs in China have been institutionalized into platform governance frameworks, actively assisting platforms in managing, monitoring, and optimizing the labor of live-streamers (Zhang and Tong, 2024; Liang and Ji, 2024; Liu et al., 2023). Unlike YouTube and other Western live-streaming platforms, where most creators remain relatively autonomous  (Alexander, 2018), Chinese live-streaming platforms rely heavily on MCNs to coordinate recruitment, training, and day-to-day performance evaluation. As a result, independent live-streamers are increasingly rare, and platform labor is mediated through MCNs that function as both infrastructural supports and informal managers (Gardner and Lehnert, 2016). These MCNs often promote their technical and analytic capacity to “manufacture” success by decoding and exploiting algorithmic mechanisms, branding themselves as algorithm-savvy “talent incubators” (Zhang and Tong, 2024; Zhang et al., 2024).
MCNs act as interpretive infrastructures in the algorithmic era. Their authority stems not from employment arrangements but from their ability to decode opaque platform systems, train streamers to meet algorithmic expectations, and define what counts as “desirable” labor. This role is especially critical in China’s live-streaming industry, where many streamers, often rural migrants or young workers with limited digital literacy (Zhao, 2024; Tang et al., 2022; Xiao et al., 2024a), rely on MCNs as their main, sometimes only, source of guidance. Zhang and Tong recently offered a systematic account of the historical development of the MCN ecosystem in China and argued that Chinese content platforms highly rely on MCNs to absorb the risks of professional content production  (Zhang and Tong, 2024). They also acknowledge that MCNs play an active role in helping creators become familiar with platform algorithms and optimize their content accordingly  (Zhang and Tong, 2024). However, they stop short of investigating a critical question: how do MCNs specifically construct and disseminate knowledge about algorithms, and why do they do so in particular ways? As a result, the role of MCNs in shaping creators’ epistemic relationship to platform algorithms as both interpreters and enforcers remains undertheorized. Our study addresses this gap by exploring how MCNs strategically mobilize informal algorithmic knowledge not only to support creators, but also to discipline and manage them.
3. Methods
In this study, we conducted extensive fieldwork primarily in Beijing and Changsha in China, focusing on MCN companies and live-streamers. From April to December 2023, we engaged in a nine-month in-depth ethnographic investigation. Our fieldwork involved observing the daily operations of three MCN companies (AB, QC, XH in our cases)—pseudonyms used to protect the anonymity of participating organizations—including how they set up live-streaming environments, provided technical and emotional support, and managed interactions among managers, technologists, and live-streamers. We also followed streamers’ broader daily routines to understand their labor conditions and platform practices in context. All observations were conducted with the consent and cooperation of both the MCN companies and the individual participants.
To supplement this ethnography, we conducted one-on-one online interviews with 24 live-streamers and 13 MCN practitioners, each lasting approximately one hour. While our immersive fieldwork was based in three MCNs, our interview sample expanded to include participants from a total of seven MCN companies (AB, QC, XH, KX, HG, RZ, XY).
This research received approval from the university’s Institutional Review Board (IRB). Throughout the study, we were transparent about our researcher roles and obtained informed consent from all participants. We also engaged in ongoing reflexive practices, including regular internal discussions among research team members to critically examine power dynamics, positionality, and potential sources of bias.
3.1. Participants: Live-Streamers (N=24), MCNs Practitioners (N=13)
We interviewed 24 live-streamers (L1–L24) active on major Chinese platforms including Laifeng, MoMo, YY, Douyin, and Huajiao. Participants ranged in age from 21 to 34 and came from diverse gender, educational, and geographic backgrounds (see Table 1). About two-thirds (n=15) were full-time streamers, while the remainder worked part-time due to unstable income. We also interviewed 13 MCN practitioners (M1–M13) from seven companies spanning different stages of organizational development, including startups, growth-stage firms, and mature MCNs (see Table 2). Their roles included executives (n=5), managers (n=6), and technicists (n=3), with one participant holding a hybrid manager-technician position. Executives made strategic decisions; managers coordinated scheduling and performance; and technicists provided technical support and training. We intentionally sampled  (Campbell et al., 2020) participants across a range of employment statuses, platform types, MCN development stages, and institutional roles to reflect the common experience of China’s live-streaming industry. By including voices from both workers and MCNs, our study design aims to surface not only individual experiences but also broader structural patterns of visibility, precarity, and control.
Table 1. Information of Live-Streamer Participants.
Name Gender Age MCN Work Tenure Education Platform Fans Employment Monthly Income
L1 Female 21 AB 8 months Undergraduate Laifeng 3k Part-time ¥1000-2000 / \$140-280
L2 Male 24 AB 6 months Master Laifeng 2k Part-time ¥1000-1500 / \$140-210
L3 Female 21 KX 3 months Undergraduate Laifeng 1k Part-time ¥3000-4000 / \$420-560
L4 Female 25 QC 3 months Master Laifeng 1k Part-time ¥2000-3000 / \$420-700
L5 Female 22 XH 7 months Senior high school Laifeng 23k Full-time ¥2000-4000 / \$280-560
L6 Female 26 XH 8 months Senior high school Laifeng 1k Full-time ¥1500-2500 / \$210-350
L7 Male 24 QC 3 months Junior college Laifeng 1k Full-time ¥1500-3000 / \$210-420
L8 Female 24 QC 9 months Junior college Laifeng 1k Part-time ¥2000 / \$280
L9 Female 18 KX 2 years Junior college Laifeng 6k Full-time ¥3500-4000 / \$490-560
L10 Female 23 XH 5 years Senior high school Laifeng 148k Full-time ¥10000-15000 / \$1400-2100
L11 Female 26 XH 4 years Undergraduate Laifeng 98k Full-time ¥10000-12000 / \$1400-1680
L12 Female 28 XH 2 years Senior high school Laifeng 57k Full-time ¥8000-10000 / \$1120-1400
L13 Female 20 XY 3 years Junior college MoMo 12k Full-time ¥3000-6000 / \$420-840
L14 Female 28 XY 1 year Senior high school MoMo 1k Full-time ¥4000-4500 / \$560-630
L15 Female 26 XY 1.5 years Junior college MoMo 1k Full-time ¥3000-5000 / \$420-700
L16 Female 24 RZ 2.5 years Junior college $Y$ 2k Full-time ¥1000-3500 / \$140-490
L17 Female 28 RZ 1 year Undergraduate $Y$ 1k Full-time ¥2000-4000 / \$240-560
L18 Female 23 XH 1 year Undergraduate Douyin 2k Full-time ¥1500-2000 / \$210-280
L19 Male 21 XH 1.5 years Undergraduate Douyin 83k Part-time ¥2000-3000 / \$280-420
L20 Male 24 RZ 2 years Undergraduate Douyin 75k Part-time ¥3000-4000 / \$420-560
L21 Male 34 HG 3 months Undergraduate Douyin 2k Part-time ¥1000-1500 / \$140-210
L22 Male 26 HG 4 months Master Douyin 9k Part-time ¥2000-2500 / \$280-350
L23 Male 29 RZ 2 years Junior college Douyin 15k Full-time ¥4000-5000 / \$560-700
L24 Female 32 RZ 1 year Master Huajiao 3k Full-time ¥3000-3500 / \$420-490
Table 2. Information of MCNs Participants
Name(ID) MCNs MCNs Stage Position Gender Age Working Seniority Education
M1 QC Mature Executive Male 42 4 years Master
M2 KX Startup Manager Female 35 2 years Junior college
M3 AB Mature Executive Female 42 5 years Undergraduate
M4 XH Growth Manager/Technicist Male 40 4 years Junior college
M5 AB Mature Manager Female 40 1 year Undergraduate
M6 XH Growth Executive Male 48 5 years Junior college
M7 RZ Startup Executive Female 46 2 years Ph.D
M8 HG Startup Manager Male 53 2 years Undergraduate
M9 XY Mature Manager Male 25 2 years Junior college
M10 RZ Startup Manager Female 35 2 years Master
M11 XY Mature Executive Male 30 3 years Junior college
M12 XH Growth Technicist Male 24 1 year Junior college
M13 XH Growth Technicist Male 22 6 months Undergraduate
3.2. Study Design
3.2.1. Nine-month Ethnography in Beijing and Changsha
We conducted a nine-month ethnographic study (April–December 2023) within three MCNs (AB, QC, XH) based in Beijing and Changsha, actively participating in their live-streaming practices. These two cities were chosen because they represent distinct yet influential hubs in China’s live-streaming economy: Beijing, as the nation’s capital, is a center of digital entrepreneurship and policy discourse (Cunningham et al., 2019); Changsha, though a smaller city by comparison, has emerged as a leading cultural production hub that attracts a large number of rural migrant streamers due to its thriving live-streaming industry (Xiao et al., 2024a).
Over the course of our nine-month ethnographic fieldwork, we focused on how MCNs and live-streamers interact around platform algorithms. In the first three months, we took an open and exploratory approach to learn how streamers get started, what their daily routines look like, and how MCNs provide support or step in to guide their work. During this period, we also picked up basic streaming skills ourselves and built trust with both streamers and MCN staff. In the middle phase of our fieldwork, we started noticing how often both MCNs and streamers talked about “the algorithm.” The algorithm wasn’t just seen as background technology but became a key part of everyday decision-making. MCNs used it to set performance expectations, give advice, and explain success or failure. Streamers, in turn, developed their own informal beliefs about how the algorithm works, often based on MCN coaching. In the final four months, we shifted our attention fully to this process. We closely examined how MCNs interpret vague and unpredictable algorithmic signals and turn them into concrete rules and routines for streamers to follow. This shift led us to understand MCNs not just as business managers, but as algorithmic intermediaries, the organizations that translate platform uncertainty into structured labor expectations through shared stories, metrics, and coaching strategies.
We believe ethnography is particularly well suited for this research because it enables access to the tacit routines, emotional investments, and situated negotiations that structure algorithmic labor, the factors often overlooked in formal accounts or interviews. Long-term immersion allowed us to observe how such folk theories of algorithms are socially constructed and organizationally reinforced, not just individually imagined.
3.2.2. Interviews with Live-Streamers and MCNs Practitioners
Semi-structured interviews are widely recognized as a valuable complement to ethnographic fieldwork, offering opportunities to explore emergent themes in greater depth (Thach et al., 2023). Following our nine-month ethnography, we conducted semi-structured interviews in the end of our ethnography with 24 live-streamers and 13 MCN practitioners to further investigate how streamers perceive algorithmic control and how MCNs interpret, mediate, or reinforce these algorithmic dynamics.
Participant recruitment began with the three MCNs involved in our fieldwork, who helped us access a broader network of streamers and affiliated MCNs, to our final sampled seven MCNs, allowing us to diversify our sample beyond the initial field sites. This expansion was necessary for two reasons. First, MCNs vary considerably in size, specialization, and degrees of platform dependency, which can lead to various algorithmic strategies and labor practices. Second, broadening the sample helped us identify cross-organizational common patterns, enabling a more robust understanding of how algorithmic knowledge is constructed and circulated across the MCN ecosystem. We intentionally included both novice and experienced streamers, as well as MCN staff from different roles and organizational levels.
The interview protocol, developed based on prior research on live-streamers (Wohn and Freeman, 2020; Wang et al., 2019; Wu et al., 2023) and MCNs (Vonderau, 2016; Liang, 2022; Siciliano, 2023), then focused on several core areas: challenges associated with platform algorithms, streamers’ understanding and navigation of algorithmic systems, the forms of algorithmic guidance or control exerted by MCNs, and the infrastructural or interpretive support MCNs provide to facilitate compliance or success. For live-streamers, sample questions included: “How do you know when the algorithm is working in your favor or against you?”; “Has your MCN ever provided you with explanations or advice about the algorithm?”; and “What kind of support—technical, emotional, or interpretive—do you receive from your MCN in navigating platform challenges?” For MCN practitioners, we asked: “How do you stay informed about changes in the platform’s algorithmic rules?”; “Do you provide training or informal guidance to streamers on how to ‘play to the algorithm’?”; and “Would you describe your role for live-streamers regarding platform algorithms?”
Each interview lasted approximately 60 to 90 minutes. Throughout, we encouraged participants to critically reflect on how platform algorithms shape their daily labor and how MCNs influence these experiences, not only as organizational actors but also as mediators of algorithmic knowledge and expectations.
3.3. Data Analysis
All interviews were conducted and recorded in Chinese to ensure participants could communicate comfortably and effectively. Three researchers subsequently transcribed and translated these recordings into English for analysis. We conducted a thematic analysis, combining both deductive and inductive coding methods. Deductive codes were informed by our theoretical framework and interview protocol, centering on constructs such as algorithmic governance, mediation, risk transfer, and infrastructural support. Inductive codes emerged through open reading of the transcripts and fieldnotes, capturing unanticipated practices, metaphors, tensions, and expressions from participants. We analyzed our ethnographic fieldnotes and interview transcripts together, rather than treating them as separate datasets. This integrated approach allowed us to triangulate across different forms of data, what participants said in interviews and what we observed in situ, yielding richer contextualization and interpretive depth  (Xiao et al., 2024b). Drawing from traditions in interpretive ethnography and qualitative HCI, we treated interviews not as isolated accounts but as situated reflections shaped by ongoing organizational and algorithmic conditions observed during fieldwork  (Seaver, 2017).
We adopted a reflexive thematic analysis approach  (Kuo et al., 2023; Xiao et al., 2024b), which foregrounds the researchers’ interpretive labor and views knowledge production as iterative and situated. To begin, all three researchers independently coded a subset of transcripts and fieldnotes to generate initial codes. These were compared in regular team discussions, where we negotiated code definitions, added emergent categories, and refined our shared codebook. Rather than calculate inter-coder reliability, we emphasized interpretive alignment through dialogue, clarification, and documentation of divergent readings, following qualitative HCI best practices. After completing the open coding phase, we clustered related codes into higher-order conceptual categories. We used affinity diagramming (Haskins Lisle et al., 2020) for organizing codes, identifying cross-cutting themes, and mapping connections between concepts. Throughout the process, we wrote analytic memos to document interpretive decisions, surprising findings, and theoretical insights, maintaining a reflexive audit trail that anchored our evolving understanding.
3.4. Positionality Statements
Our research team brings diverse yet complementary perspectives shaped by years of engagement with the live-streaming ecosystem. Five researchers in our team have been involved in long-term research experience into platform labor and digital media economies. One team member has conducted extensive collaborative work with MCN companies, affording us privileged access to organizational practices and streamers’ networks while also requiring ongoing reflexivity about the power dynamics embedded in such relationships. Another two researchers are trained in critical political economy and labor studies, providing a structural and historically grounded lens to interpret data related to labor precarity. One member has conducted multiple ethnographic and interview-based studies on live-streaming cultures and creators’ self-presentation, contributing cultural and interactional sensitivity to our analysis. This range of backgrounds enriched our analytical process while also necessitating deliberate reflection on how our disciplinary commitments and field relationships shaped the way we framed questions, interpreted participants’ accounts, and conceptualized key themes.
4. Findings: Constructing Dual Folk Theories: How MCNs Make Opaque Algorithms Actionable for Streamers? (RQ1)
Our findings show that both MCNs and live-streamers operate under the same fundamental condition: algorithmic opacity. Live-streaming platforms rarely disclose how their recommendation systems function. As a result, both MCNs and streamers must navigate a system that is unpredictable, constantly shifting, and difficult to decipher. For streamers, this opacity often leads to frustration and powerlessness. They cannot understand why one stream gains traction while another goes unnoticed, or why their platform visibility drops without warning. In search of guidance, many turn to MCNs and hope these institutions have insight into how to improve their algorithmic exposure. However, MCNs themselves are not privileged insiders. Like the streamers they manage, they do not have access to the algorithm’s inner workings. They are also left guessing.
To make sense of this uncertainty, MCNs observe performance trends, compare outcomes across multiple streamers, and develop what we call folk algorithmic theories: informal, experience-based beliefs about how the algorithm behaves, what kinds of content it rewards, and how it might be influenced. These folk beliefs are speculative rather than technical. They are grounded in trial-and-error and pattern recognition, not in verified mechanisms. As many MCN interviewees in our study admitted, there is no guaranteed formula for algorithmic success, no matter how closely streamers follow the advice they are given.
In other words, MCNs are fully aware that their guidance cannot ensure that any individual streamer will succeed under the algorithm’s logic. As a result, MCNs maintain two distinct but interconnected versions of algorithmic theory. Internally, MCNs recognize that the algorithm is unstable, opaque, and ultimately uncontrollable. This understanding allows them to make business decisions cautiously, using flexible planning and risk management. Externally, however, they offer streamers simplified and more confident theories of algorithmic visibility. They portray the algorithm as fair, learnable, and responsive to effort. The message is clear: “If you keep trying, and follow our advice, you’ll succeed.” This theory of algorithms helps recruit and retain streamers, many of whom are willing to enter and stay in the industry, invest financially in equipment and services, and devote emotional energy in hopes of eventually gaining algorithmic visibility.
In the sections that follow, we show how MCNs carefully maintain these two theories—one inward-facing and probabilistic (subsection 4.1), the other outward-facing and prescriptive (subsection 4.2). Together, these two versions of algorithms theories allow MCNs to operate effectively in a deeply uncertain environment while preserving organizational authority and extracting economic value from precarious digital labor.
4.1. Internal Folk Theories: Managing Uncertainty and Risk within MCNs
Despite their professional positioning and close partnership with platforms, MCNs do not possess privileged access to the internal workings of recommendation algorithms. For many MCN staff, the algorithm remains an unpredictable and opaque black box. “We don’t have backend access,” explained an operations manager (M3). “All our understanding of the algorithm is based on repeated trial and error (M10).” Faced with this uncertainty, MCNs develop their own internally shared folk theories within company staff to interpret platform behavior and improve streamers’ algorithmic visibility.
These internal folk theories are expressed through heuristic rules and behavioral scripts, often referred to within the organizations as “formulas” or “track logics.” Typical examples include recommendations on streaming frequency (e.g., “at least three sessions per week”), minimum broadcast duration (e.g., “streams under 40 minutes won’t get recommended”) and emotional tone (e.g., “you must be excited in the first five minutes, or the system will judge it as low-interaction”). These behavioral prescriptions are formalized in onboarding materials, reiterated in weekly coaching sessions, and continuously updated via internal chat groups. Beyond behavioral scripts, many MCNs also develop aesthetic and infrastructural prescriptions that they believe enhance a streamer’s algorithmic visibility. For instance, staff often emphasize the algorithm’s supposed preference for high-definition visuals and stable audio quality, as well as the physical environment of streaming.
However, while these theories offer a sense of control, most MCN staff readily admit their limitations. “Even if you follow all the best practices, sometimes the platform just doesn’t give you any algorithmic visibility,” M9 remarked. “And we don’t really know why.” This tension between the promise of algorithmic optimization and the reality of its unpredictability has prompted many MCNs to shift focus. Rather than attempting to fully decode the algorithm, MCNs increasingly prioritize monetization logic, developing what might be called revenue-first strategies. M9 explained, ”We’ve had many cases where someone did everything right—good lighting, good engagement, great voice—and still, no algorithmic visibility. There’s just no guarantee.”
In response, MCNs adopt a statistical approach to managing uncertainty. Success is treated as a probability game: by recruiting large cohorts of new streamers and investing minimal resources in each, MCNs reduce the risk associated with any one individual’s failure. The goal is not to ensure the success of every recruit but to ensure that some will be algorithmically “picked.” As M9 elaborated: ”This process is highly random. We recruit newcomers who look attractive, provide them with some equipment, and ask them to start live-streaming. But we don’t know who will make it. What we do know is that if we have enough people, a certain percentage will succeed. Then we allocate resources to those who are ‘chosen’ by the algorithm.” This batch recruitment model allows MCNs to scale labor while minimizing upfront investment and maximizing return on attention. But the logic does not stop at internal risk distribution. Just as MCNs construct folk theories internally to manage uncertainty, they also construct simplified algorithmic theories externally—targeted at streamers.
Unlike internal folk theories, which circulate among MCN staff as tentative, probabilistic tools for interpreting platform behavior, external folk theories are simplified and confident theories constructed for streamers. These externally projected theories serve two strategic purposes: first, they attract and retain labor by portraying algorithmic success as predictable and achievable, thereby encouraging newcomers to join and existing streamers to continue; second, they stimulate financial investment by persuading streamers to purchase services, equipment, or training packages under the belief that these will enhance their algorithmic performance, allowing MCNs to profit from these sales.
While most algorithmic interpretations remain speculative for MCNs, certain practices appear to yield consistent, if not guaranteed, results. For example, some MCNs staffs (Such as M3, M4, M8, M9) report success in boosting algorithmic visibility by inputting large numbers of real-user accounts as followers for new streamers or by orchestrating coordinated entry of verified user accounts into live-stream rooms. These strategies are believed to trigger the algorithm’s visibility mechanisms, resulting in increased exposure and higher likelihood of recommendation. However, it is crucial to emphasize that even these seemingly “effective” tactics operate within a system of high variability and limited transparency. What works under one live-streaming platform version or for one streamer may not hold for another, reinforcing the broader uncertainty that underpins MCNs’ dual theorization of algorithms. Moreover, for most ordinary streamers, MCNs are often unwilling to invest significant resources into these tactics, given the high costs and uncertain returns. Thus, in this study, we do not seek to examine how MCNs develop sophisticated understandings of platform algorithms in order to manipulate them—indeed, in our fieldwork, MCNs rarely invested substantial effort in decoding or directly manipulating the algorithm, which they largely viewed as infeasible, unrewarding, and not worth the effort given the high uncertainty and low return. Rather, our focus lies in how MCNs construct and mobilize algorithmic folk theories—as organizational narratives—to manage algorithmic uncertainty, structure labor practices, and sustain control over streamers within the live-streaming ecosystem, a strategy that, in the context of our sampled Chinese MCNs, represents a more direct and prevalent path to profitability than algorithmic manipulation.
Ultimately, what finally emerges is a dual system of epistemic practice about algorithmic folk theories: internally, MCNs recognize the algorithm as volatile and uncontrollable; externally, they reframe it as predictable and improvable. This asymmetry is not incidental—it is the foundation of the MCN business model. By selectively mobilizing folk theories in different contexts, MCNs manage precarity, sustain labor supply, and extend their authority within China’s live-streaming industry. Next, in subsection 4.2, we further analyze how these externally directed folk theories of algorithms are strategically constructed by MCNs for streamers.
4.2. External Folk Theories: Simplified Narrative to Manage and Govern Streamers
In this section, we analyze how MCNs construct and disseminate folk theories of algorithms to live-streamers. Unlike the probabilistic and speculative nature of internal folk theories held by MCN staff, these outward-facing theories of algorithms targeting streamers are packaged as confident, simplified rules that make the algorithm seem legible and controllable. We identify two primary ways in which these theories are strategically deployed: (1) by translating algorithmic logic into concrete, observable metrics such as appearance and equipment quality (subsubsection 4.2.1); (2) by framing algorithmic success as attainable through discipline, investment, and compliance (subsubsection 4.2.2). These theories do not merely inform behavior that shape streamers’ beliefs, extract their investment, and sustain a system in which hope becomes both affective infrastructure and a mechanism of labor control.
4.2.1. Simplifying Algorithmic Logic Through Observable Metrics
In the face of algorithmic opacity and instability, MCNs firstly offer simplified but confident explanations to live-streamers that translate the platform’s black-box logic into concrete, observable metrics. Without access to the actual workings of recommendation systems, MCNs construct a seemingly actionable framework for live-stramers grounded in visible variables, such as physical appearance, equipment quality, and early user engagement, to make the algorithm appear understandable and responsive to effort.
For example, M2 usually explains to new streamers that physical appearance plays a decisive role in gaining algorithmic exposure by increasing early viewer retention: “The audience will only give a live-streamer three seconds to show themselves and keep them. When a live-streamer is particularly attractive or has a humorous appeal, they are more likely to gain audience favor. This results in higher viewer retention and influx rates, leading the recommendation algorithm to classify them as popular live-streamers and, consequently, grant them more exposure.” This logic, while unverified, is widely internalized by streamers.
Another common metric emphasized by MCNs is the quality of equipment. M4 explained that many platforms set technical thresholds for stream quality, and that insufficient equipment could lead to streams being algorithmically filtered out: “If your sound and visuals are poor, people drop out quickly, and that hurts your retention rate. Algorithms pick up on that. So we always recommend at least the basics, e.g., lights, microphones, stable connection.” Streamers echoed this perceived necessity. L5 noted that even renting equipment was necessary for basic visibility:“Whether renting or buying equipment, it at least gives you a chance to gain some recommended followers and earn a bit of money. Without investing in equipment, it’s nearly impossible to be visible to viewers.” As shown in Figure 1, MCNs distribute tiered technical audio equipment in ways that symbolically link production quality with algorithmic visibility, reinforcing an externally imposed folk theory that equates financial investment in equipment with success.
Figure 1. Tiered technical audio solutions deployed by MCN agencies serve as material markers of professional status and algorithmic worth within the live-streaming economy. (a) Entry-level handheld microphone provided as standard equipment for contracted streamers, a counterfeit version of the Blue E300, with production costs ranging from 60-70 RMB (approximately 8 USD). (b) Mid-tier condenser microphone offered complimentary with streaming studios, including free replacement service for malfunctions; this unit replicates the Neumann TLM 193 with manufacturing costs between 300-400 RMB (approximately 50 USD). (c) Shure SM7B microphone recommended by MCN agencies for talk show hosts, voice actors, and vocal performers with established audience bases; streamers are encouraged to purchase this equipment through the MCN at a one-time fee of 4,000 RMB (market value approximately 2,500 RMB), inclusive of complementary mixing console and technical support services. (d) Flagship audio broadcasting system typically reserved for contracted knowledge-based/educational teams or professional singers; streamers may temporarily rent this equipment for cover or original song recordings at approximately 300 RMB (40 USD) per hour, with post-production and technical assistance available at additional cost. This premium equipment primarily serves as a demonstration of the MCN’s professional capabilities and is not available for external purchase.
What often begins as a recommendation to improve a single technical element, such as audio clarity or lighting, gradually expands into a broader vision of the “ideal” streaming setup. MCNs do not simply promote individual tools; they encourage streamers to view the entire streaming environment as a form of algorithmic currency. Ring lights, DSLR cameras, soundproofed backdrops, multi-angle mounts, branded filters, and LED mood lighting are framed as signs of professionalism and algorithm-readiness. As streamers internalize this logic, the line between technical sufficiency and aesthetic perfection becomes blurred. A basic setup is no longer seen as enough to ”meet the algorithm’s baseline”; instead, streamers feel pressured to constantly upgrade in order to stay competitive within an ever-rising standard.
However, most MCN staff themselves acknowledged the limitations of this external theory for streamers (which they don’t believe internally). M9 pointed out the asymmetry between perceived causes of success and actual algorithmic outcomes: “We can assume that average looks or limited equipment make it difficult to receive algorithmic recommendations. However, this assumption cannot be reversed. We can’t conclude that being attractive or having better equipment will guarantee algorithmic recommendations. Our company’s data doesn’t support this.”
Despite this internal skepticism, MCNs continue to promote the idea to live-streamers that algorithmic success is predictable and dependent on specific surface-level traits. They construct the algorithm as a controllable input-output system: if streamers follow the right steps, if maintain a certain visual style, upgrade equipment, stream for long enough, and use the right tags, they are told they will be rewarded with visibility and growth. This framing simplifies algorithmic complexity into a checklist of behaviors and attributes, turning uncertainty into a set of seemingly manageable variables.
Because this theory of algorithm constructed by MCNs feels actionable, most streamers often accept it without question. They internalize the idea that if they fail to gain traction, it is because one of these variables was insufficient. Perhaps their lighting was not professional enough, their background not aesthetic enough, their streaming duration too short, or their performance too flat. This belief system allows MCNs to continuously raise performance benchmarks, always pointing to another metric that could be optimized. For streamers, each failed attempt becomes a reason to try harder, upgrade further, or self-correct, rather than question the underlying opacity or instability of the algorithm itself. In this way, MCNs maintain authority while pushing laborers into a cycle of self-blame and incremental reinvestment. Crucially, this theory also creates a steady revenue stream for MCNs, as streamers are encouraged to repeatedly invest—purchasing equipment, renting better spaces, or enrolling in paid training sessions—all under the belief that these expenditures will enhance their algorithmic visibility and lead to success.
4.2.2. Framing Algorithmic Success as Attainable and Controllable
A second common way MCNs explain the algorithm is by framing it as attainable and controllable, a fair system that rewards effort, discipline, and the right technical setup. Despite the algorithm’s underlying opacity and volatility, MCNs portray it as a rational mechanism that can be “understood” and “tamed” through compliance with certain behavioral and technical standards. This framing offers streamers a sense of structure and hope, suggesting that visibility is within reach if they follow instructions and persist.
L3 described how this theory of algorithms shaped her initial investment in equipment: “Those who can’t afford live-streaming equipment have to rely on income from other sources to make an initial investment for the equipment [if they believe live-streaming could earn more money].”
This theory is not just implicit. It is actively constructed by MCNs during recruitment and onboarding. New streamers are told that the algorithm can be “understood” and “tamed,” and that their success depends on how closely they adhere to MCN advice and technical standards. This framing channels attention away from platform bias or structural inequality, and toward individual action and responsibility.
Crucially, MCNs often present live-streaming platforms and their algorithmic infrastructures as less discriminatory and more open than other forms of labor or public-facing institutions. In particular, our sampled MCNs in China emphasize that streamers from rural, migrant, or female backgrounds may find more opportunities online than in traditional workplaces. Unlike offline industries where class, gender, accent, or education may limit mobility, MCNs describe algorithmic systems as blind to such characteristics, privileging performance, persistence, and technique instead. Several streamer interviewees (n=7) recalled being told during onboarding sessions that “the algorithm doesn’t care who you are, only what you do”, a message that resonated deeply with streamers from marginalized backgrounds.
This perception helps explain why live-streaming, despite its precarity, continues to attract disproportionately high numbers of rural migrants and young women. The promise of algorithmic neutrality, reinforced by MCNs’ messaging, offers both a psychological appeal and a perceived escape from entrenched discrimination elsewhere. By portraying the algorithm as neutral and open to anyone, MCNs produce not just training routines for live-streamers, but an aspirational and moral framework in which effort is rewarded, and structural inequalities appear bypassed.
However, as most of MCNs interviewees (N=8) agreed, the actual support that MCNs provide is often minimal. They may offer basic resources, such as startup equipment, help with platform registration, and initial streaming advice, but rarely offer deeper guidance about how to gain algorithmic visibility. What matters more is the theories of algorithms they construct: that the algorithm is fair, that it can be tamed, and that success is attainable through individual effort. This story is powerful not because it is empirically accurate, but because it motivates.
For MCNs, this external version about theory of algorithms serves a clear purpose: it attracts a continuous influx of new streamers. Even if only a small percentage achieve visibility, the system is sustained by the hope of success and the belief that the algorithm rewards discipline. For streamers, especially those from marginalized backgrounds with few algorithm literacy, this external folk theory provides a rare sense of possibility, even if the odds remain opaque and the outcomes unpredictable.
Crucially, it is precisely this belief that the algorithm can be learned, tamed, and ultimately trusted sustains repeated investment from streamers. Streamers are told that success is not arbitrary, but a matter of strategy, effort, and consistency. The algorithm, unlike human employers or traditional gatekeepers, is portrayed as objective and responsive by MCNs. For many streamers, especially those excluded from stable or upwardly mobile labor markets, the promise of algorithmic fairness feels both empowering and worth pursuing. This perception justifies the continued purchase of equipment, the adoption of performance routines, and the sacrifice of time and emotional energy. Even when results fall short, the theories remains intact: if one stream fails, the next might succeed. In this way, MCNs not only mobilize belief—they build a system in which belief becomes labor, and hope becomes infrastructure.
5. Findings: Institutionalizing Folk Theories: How MCNs Operationalize Algorithmic Beliefs for Labor Management (RQ2)
Building on section 4, which examined how MCNs construct internal and external folk theories of algorithms, we now turn to how these interpretations are institutionalized into labor control. First, MCNs embed algorithmic expectations into everyday routines, through onboarding, coaching, and evaluations, repeating the message that effort, emotion, and consistency will be rewarded by the algorithm. This moralized framing aligns with the external folk theory that success is earned, motivating streamers to self-discipline and persist. At the same time, MCNs internally acknowledge the algorithm’s volatility and shift the burden of failure onto streamers through this moral frame, who blame themselves rather than systemic opacity (subsection 5.1). Second, MCNs formalize these beliefs through structured schedules and aesthetic standards, reinforcing the idea that algorithmic success can be optimized. Streamers are encouraged to upgrade equipment, refine routines, and purchase services—all justified by the external folk theory. Yet MCNs, guided by internal theories, understand the algorithm offers no guarantees. Their profits rely not on streamer success, but on sustaining belief and dependence on paid services (subsection 5.2).
5.1. Transforming Platform Speculation into Moral Labor Expectations
As discussed in the previous subsection 4.2, MCNs interpret opaque and unstable platform algorithms by translating them into simplified rules and actionable advice. These interpretations are not neutral; they serve specific organizational interests by creating the illusion that algorithmic success can be achieved through discipline, effort, and optimization. In this section, we show how such interpretations do not simply circulate as loose theories. They are systematically leveraged to shape how streamers work, behave, and evaluate themselves.
One major way this happens is through the transformation of speculative ideas about the algorithm into moral labor expectations. Beliefs such as “the algorithm rewards long hours” or “visual quality guarantees recommendations” are framed not as possibilities, but as certainties. These claims are repeatedly emphasized in onboarding materials, group chat guidance, and informal coaching, where they gain authority as commonsense truths. Over time, they become internalized by streamers as standards of professional conduct.
This process marks the point where algorithmic interpretation becomes algorithmic governance. Instead of merely helping streamers understand how the platform might work, MCNs use these beliefs to define how streamers should behave. The burden of visibility is shifted entirely onto the individual: if streamers are not being recommended, it is because they are not working hard enough, not investing enough, or not attractive or expressive enough. L6 recalled: “The MCNs I contacted told me I only needed to stream every day. Just show up, and the algorithm will recognize you. I thought, okay, then it’s just about effort.”
This framing naturalizes structural opacity as personal failure. M2 explained: “Most streamers think it’s their fault. That’s what keeps them going. They keep chasing that moment when the algorithm ‘notices’ them.” Even when outcomes do not improve, many streamers continue to blame themselves. L7 described her own response: “They said if I streamed every day and did more facial expressions, I’d get algorithmic recommendations. So I smile even when I’m exhausted, because maybe today is the day the algorithm sees me.”
These internalized expectations stem directly from the way MCNs frame the algorithm in the first place for streamers, as something knowable, responsive, and fair. Because streamers believe the system rewards the right behavior, they are willing to endure exhaustion, self-doubt, and financial risk. This makes the MCN’s job easier. As M11 put it: “We MCNs don’t need to spend much. If they streamers believe in the algorithm, they’ll work hard for free. That’s the best-case scenario for any company.”
Streamers are not being forced to work harder; they are being persuaded that working harder is their own path to success. MCNs reinforce this belief system by presenting guidance not as command, but as mentorship. L18 explained: “My manager never says I must do something. She just says, ‘Other streamers who do X or Y tend to get recommended more.’ Then she might show me screenshots. It’s never a command, but I always feel like if I don’t follow, I’ll miss out.”
What appears as care or mentorship is, in fact, a form of indirect control. Streamers act based on what they believe the algorithm expects, but those beliefs are shaped by the MCN’s theories. In L22’s case, her makeup choices were labeled as insufficient by MCNs not because of any clear platform rule, but because the MCN had already instilled the idea that visual performance is central to algorithmic success. This claim itself emerges from the broader uncertainty of algorithmic systems: in the absence of transparent criteria, MCNs gain interpretive authority, turning speculation into behavioral norms. As L22 put it, “They didn’t force me, but I felt like if I didn’t improve my makeup, I was being lazy or unprofessional. It felt like a personal failing.”
These internalized expectations not only sustain streamers’ labor but also strategically align with the MCNs’ internal understanding of algorithmic uncertainty. Internally, MCNs know they cannot guarantee any single streamer’s success due to the volatility of recommendation algorithms. As such, offloading the responsibility of failure becomes an essential managerial tactic. It allows MCNs to avoid accountability for poor performance outcomes while maintaining the appearance of support and guidance.
5.2. Embedding Algorithmic Beliefs into Material and Contractual Structures
The interpretive work described in the previous section does not stop at narrative or motivation. MCNs extend their algorithmic explanations into the material, aesthetic, and contractual arrangements that define how streamers work. What begins as speculative beliefs, such as the idea that visual clarity or consistent appearance leads to higher algorithmic visibility, is gradually institutionalized into binding routines and financial commitments.
A central mechanism is the pressured acquisition of high-end equipment. MCNs frequently tell streamers that clearer visuals and smoother audio will improve their ranking, reinforcing this belief through formal or informal expectations. L19 described her experience: “I spent 10,000 RMB on a Sony camera because the MCN said it would help me algorithmic visibility rank better. But nothing changed. They kept saying, maybe I need to improve my makeup or get a new background.”
This expectation extends far beyond cameras. Ring lights, microphones, mixers, reflective backdrops, and LED panels are all framed as essential to “meeting the algorithm’s baseline.” In some cases, MCNs distribute internal “technical evaluation reports” that nudge streamers toward upgrades. M12 explained how this theory benefits the company: “Most streamers don’t buy equipment all at once. They rent from us first. If they start earning, they’ll invest more. If they don’t, at least we make money from the rentals.”
Alongside hardware, MCNs standardize streamers’ aesthetic presentation. Guidelines often dictate not only tone and posture, but makeup, camera angles, and even clothing schedules. M9 described a typical protocol: “They’re not allowed to wear the same outfit in a two-week cycle. We even advise which colors look better on camera. Some of them rent clothes from us. We keep it on their account.”
These requirements are enforced not just rhetorically, but contractually through studio rules, performance evaluations, and surveillance systems. Even streamers who work from home are monitored. M4 explained: “We ask them to share their live link with a manager who checks if the lighting, camera, background are aligned with expectations. If not, they get a warning. If the setup is too basic, we suggest they pause streaming until they upgrade.”
In some cases, MCNs also manage the physical layout of labor. Many operate shared studio spaces where streaming rooms are tiered by performance level. Higher-ranked streamers receive better infrastructure and services. L16 described how space and performance were linked: “At first, I shared a basement room with three other girls. It was tiny, and the internet cut out often. When I ranked higher one month, I moved to a private room upstairs. But I had to pay more for it, nearly 3,000 RMB a month.” These layered expectations create cycles of escalating cost and performance pressure. L20 recalled: “I spent most of my early earnings on room rental, clothes, gear. I didn’t even realize I was in debt until month three. But by then, I didn’t dare quit; I was too invested.” Streamers are often told that these are necessary “investments” in their future algorithmic visibility. But the platform algorithm offers no guarantees. L22 shared her experience: “I borrowed money from my cousin to redecorate my stream room. I told myself, if I looked more professional, I’d get better exposure. But nothing happened. Now I just feel stuck.” As shown in  Figure 2, the material differentiation of studio spaces reflects a managed hierarchy of visibility and professionalism, reinforcing MCNs’ externally imposed belief that algorithmic success is contingent on aesthetic and technical upgrades.
Figure 2. Comparative analysis of livestreaming studios with varying amenities and price points. All studios provide basic air conditioning, internet connectivity, and workspace furnishings, while differing substantially in equipment quality and decorative elements. (a) A bare basement dance streaming studio featuring a faux window facing a concrete wall, priced at approximately 2,500 RMB monthly without computer equipment, lighting systems, or professional backdrops beyond a rudimentary fringe curtain. (b) A compact basement studio designed for singing and conversational streaming, characterized by its confined windowless space, basic interior finishing, and pink swivel chair, available at approximately 2,200 RMB monthly. (c) A ground-floor multipurpose streaming studio equipped with fundamental lighting apparatus and computer systems, featuring a dedicated window, priced at approximately 3,100 RMB monthly. (d) A premium spacious studio situated on the second floor, furnished with an independent clothing rack, basic lighting equipment, professional microphone with stand, in-room audio monitoring speakers, dedicated window access, and a vanity mirror, available at 3,500 RMB monthly with the additional advantage of proximity to technical support staff offices for immediate assistance.
As illustrated in subsubsection 4.2.1, MCNs build streamers’ belief that algorithmic success depends on visible performance cues and technical polish. Once internalized, these beliefs become a foundation for control: they justify expenditures, reinforce discipline, and make streamers responsible for conditions over which they have limited influence. M12 captured the disciplining effect of this folk algorithmic theory: “We don’t force anyone to stay. But when you’ve rented equipment, decorated your room, signed the platform contract; it’s hard to walk away. You want to try a bit longer, just in case.”
For MCNs, this internalization is not just a behavioral outcome—it is a business model. Once streamers adopt the belief that algorithmic success can be engineered through effort and appearance, they become willing customers for MCN-provided goods and services: studio rentals, training sessions, lighting kits, camera upgrades, and branding packages. These purchases are not simply one-time decisions, they form an ongoing cycle of reinvestment, driven by the hope that one more improvement will finally unlock algorithmic visibility.
This behavior logic aligns precisely with MCNs’ internal understanding of algorithmic unpredictability. Internally, MCNs acknowledge that algorithmic success cannot be reliably manufactured at the individual level. As such, they prefer to spread risk across a broad base of streamers, investing minimally while maximizing returns through volume and service monetization. The more streamers believe in the narrative that effort yields exposure, the more they will spend, regardless of outcome. In this way, the externally constructed folk theory (that visibility is earned through discipline and polish) serves the internal folk theory (that success is probabilistic and unpredictable) by ensuring that MCNs profit from both success and failure, so long as belief is sustained.
It is also important to note that MCNs’ business model depends heavily on a continuous influx of new streamers who are willing to believe in the algorithmic theories MCNs constructs, as well as on retaining those who persist in pursuit of algorithmic success. For MCNs, profitability is sustained not by the success of every individual streamer, but by the broader system’s ability to motivate belief and continued investment. As such, streamers who reject these curated and institutionalized algorithmic theories or choose to exit the system early are largely inconsequential to MCNs’ operational logic. What matters is not universal buy-in, but a sufficient flow of believers willing to participate in the speculative game of algorithmic visibility.
6. Discussion
This study reveals how folk theories of algorithms—informal beliefs about how algorithmic systems work—are not merely grassroots knowledge among digital workers, but are actively constructed, formalized, and exploited by intermediary institutions. Drawing on nine months of ethnographic fieldwork in China’s live-streaming industry, we show how Multi-Channel Networks (MCNs) strategically develop dual theories about algorithms: internally, they treat algorithms as opaque and unpredictable; externally, they frame algorithms as learnable and responsive. These theories do not simply interpret the algorithm—they organize labor around it. These dual theories of algorithms, what we call internal and external folk theories, form the epistemic foundation upon which MCNs build a regime of algorithmic labor control.
This perspective complements prior live-streaming research in a critical approach. Prior studies have highlighted how live-streaming offers new forms of economic participation, cultural visibility, and affective connection, particularly for marginalized groups such as rural youth, migrant workers, and women (Freeman and Wohn, 2020; Tang et al., 2022; Momin, 2023). MCNs, in turn, have been credited with providing technical support, professional development, and algorithmic literacy, helping new entrants navigate complex platform infrastructures (Zhang and Tong, 2024; Liu et al., 2023; Zhang et al., 2024; Xiao et al., 2024a). We do not dispute these contributions. Indeed, many of our participants expressed appreciation for the opportunities afforded by both platforms and MCNs. For streamers with limited educational backgrounds, rural origins, or unstable offline employment histories, live-streaming was described as one of the few accessible avenues to economic independence and social recognition (Xiao et al., 2024a). MCNs were often credited with lowering entry barriers providing training, equipment, and algorithmic guidance that participants felt they could not have accessed on their own. Yet, our study foregrounds the structural dynamics beneath this perceived empowerment: how organizationally curated algorithmic knowledge disciplines labor without assuming direct managerial responsibility.
Our findings suggest that algorithmic folk theories function not merely as interpretive tools, but as organizational infrastructures through which labor is recruited, directed, and managed. While much of the CSCW literature has emphasized folk theories as emergent, user-driven responses to algorithmic opacity  (Shen et al., 2021; Li et al., 2023; Deng et al., 2023a), our study shifts the analytic focus to institutional actors and their role in constructing and instrumentalizing such theories. To develop this argument, the discussion proceeds in three parts. First, we revisit the concept of folk theories of algorithms, showing how they move beyond grassroots meaning-making to become institutionally manufactured knowledge regimes. We position this as a shift from individual or community-based interpretation to organizational construction (subsection 6.1). Second, we analyze how these institutionally constructed theories serve as instruments of labor governance, enabling MCNs to manage precarity, displace responsibility, and stabilize labor output through soft control (subsection 6.2). Third, we reflect on the broader design and policy implications of these findings. We call for systems and policies that support epistemic transparency, making visible not just what algorithms do, but how knowledge about them is constructed (subsection 6.3).
6.1. From Individual / Community Interpretation of Algorithms to Institutional Construction
A key contribution of this study is to reposition folk theories of algorithms not simply as informal beliefs formed through individual experience, but as forms of knowledge that can be constructed, circulated, and operationalized by organizations for specific managerial and strategic ends. Much of the existing scholarship on folk theories of algorithms has conceptualized them as informal, bottom-up interpretive frameworks that users or communities develop to make sense of opaque algorithmic systems (Bucher, 2018; DeVito et al., 2017; Eslami et al., 2016; Xiao et al., 2024b). These theories are often described as experiential, situated, and emergent, rooted in users’ lived interactions with platforms and shaped through trial and error, peer discussion, and rumor. In this framing, folk theories appear as adaptive tools, helping users navigate digital environments despite uncertainty and asymmetry in platform knowledge. Our study challenges this grassroots-centric view by demonstrating how folk algorithmic knowledge can also be systematically constructed, curated, and institutionalized by organizations. Rather than emerging only from individual experience, folk theories in the context of MCNs are strategically produced and maintained as part of the organization’s epistemic infrastructure. They are stabilized through onboarding routines, coaching scripts, and training materials; circulated via internal handbooks and chats; and selectively modified in response to shifts in algorithmic outcomes. In this process, the algorithm is rendered legible not through spontaneous interpretation, but through deliberate institutional algorithmic theory.
This shift from individuals or communities interpreting algorithmic behavior on their own, to institutions like MCNs actively constructing and distributing standardized interpretations carries several important theoretical implications for CSCW. First, we argue that folk algorithmic knowledge should be reconceptualized as infrastructural, rather than incidental. Previous literature often treats folk theories as emergent, ad-hoc, and peripheral, arising spontaneously when users try to make sense of black-boxed systems. However, our findings show that intermediary organizations such as MCNs deliberately assemble, stabilize, and disseminate particular versions of these theories. They do so not as casual helpers, but as epistemic intermediaries—institutions that claim the authority to explain what the algorithm does, how it works, and what streamers need to do to succeed. In this sense, folk theories of algorithms are not just interpretive stories, but part of the working infrastructure (Pipek and Wulf, 2009; Star and Ruhleder, 1996; Karasti and Blomberg, 2018; Ludwig et al., 2018) that governs digital labor.
Second, our findings challenge the prevailing binary in CSCW and HCI that separates algorithmic knowledge into formal (e.g., platform documentation, API guides) and informal (e.g., user speculation, peer discussion, online rumor) categories (Kusk and Nouwens, 2022; Xiao et al., 2024b; Bishop, 2019). In the context of MCNs, this division collapses. The algorithmic knowledge that circulates within these organizations is neither officially sanctioned nor casually improvised. Instead, we conceptualize it as institutionalized folk theory—informal knowledge that is produced and stabilized by intermediary actors who lack direct access to platform-engineered formal knowledge. MCNs are not platform insiders; they operate without formal connections to algorithm developers or privileged access to system logics. Yet they construct and disseminate their own working theories of “how the algorithm works” and embed these into onboarding scripts, coaching sessions, ranking systems, and contractual expectations. These folk theories gain a quasi-formal status—not through technical verification, but through routinized organizational practice, repetition, and discursive reinforcement. In doing so, MCNs transform folk sensemaking into a mechanism of labor governance.
Third, our findings broaden CSCW’s understanding of how algorithmic knowledge circulates in everyday life. Prior work has shown that workers and users build collective knowledge through grassroots means like online forums, creator communities, or worker-led organizing (DeVito et al., 2017; Karizat et al., 2021; Xiao et al., 2024b). These are important, but they are not the only sites where algorithmic knowledge is produced. Our study shows that intermediary institutions like MCNs also serve as powerful knowledge brokers. They don’t just transmit information but also shape and curate it. They decide which explanations of the algorithm get repeated, which practices get standardized, and which beliefs get embedded into training and labor management. Unlike peer-to-peer communities, MCNs combine technical interpretation with organizational authority. They have the resources and legitimacy to institutionalize speculative beliefs into widely accepted norms. In doing so, they turn uncertainty into order, and speculation into action, making algorithmic systems workable, even if not fully knowable.
Our study thus pushes CSCW to look beyond individuals or user groups when studying how people interact with algorithmic systems. It invites scholars to future consider the organizational infrastructures that mediate algorithmic knowledge and how those infrastructures shape not only what workers believe, but how they work, comply, and sustain hope.
6.2. Folk Theories of Algorithms for Labor Management
While most scholarship on folk theories emphasized their adaptive or empowering qualities, helping users make sense of opaque systems in the absence of formal knowledge (Huang et al., 2022; Siles et al., 2020; Bonini and Treré, 2024; Eslami et al., 2016), our findings suggest a more complex and strategic function: folk theories are not only interpretive frameworks, but also governance tools. In the context of live-streaming platforms, they become a key mechanism through which institutions like MCNs direct labor, manage uncertainty, and externalize risk. This repositions folk algorithmic theories from being user-centric artifacts to organizational instruments of soft control.
A growing body of CSCW and HCI literature has investigated the dynamics of algorithmic management, especially in gig economies and platform labor systems (Kellogg et al., 2020; Möhlmann et al., 2021). Here, scholars have shown how algorithms allocate work, evaluate performance, and discipline workers without the need for traditional human supervision. However, this line of work often focuses on direct algorithmic feedback (e.g., ratings, scores, rankings) or automated decision-making (e.g., job assignments). Our findings build on this by revealing a second layer of algorithmic management: the interpretive layer, where algorithmic outputs are not simply accepted or resisted, but actively narrated, framed, and naturalized through folk theories constructed by intermediaries. These folk theories are not mere byproducts of algorithmic opacity but are engineered and circulated by MCNs to align with institutional interests and govern labor at scale.
Given the opacity of platform algorithms  (Ma et al., 2025; Deng et al., 2023b; Nagaraj Rao et al., 2025), most workers lack access to how algorithmic decisions are made. While often seen as a limitation, in our case, opacity can also be leveraged as strategic ambiguity  (Gillespie, 2014; Ananny and Crawford, 2018; Eubanks, 2018). Our study shows how MCNs capitalize on this: in the absence of verified knowledge, they develop external simplified theories of algorithmic causality for streamers to set expectations, justify pressure, and shape behavior, without guaranteeing outcomes.
These external folk theories function as proxies for algorithmic authority. They tell streamers what the algorithm “wants” and how to “win” it over, but the logic they offer is shaped by MCNs’ managerial goals. By promoting these ideas as common sense, MCNs shift responsibility for failure onto streamers. Poor outcomes are framed as personal shortcomings, such as insufficient effort, investment, or affect, rather than algorithmic opacity.
Notably, this mode of governance does not depend on technical accuracy. As our findings show, even MCN staff admit their advice is speculative. But as long as these external theories appear confident, actionable and correlate with visible metrics, they retain influence to streamers. This performativity echoes insights from critical algorithm studies  (Gillespie, 2014; Eubanks, 2018): sometimes belief in algorithmic logic shapes behavior, regardless of its truth.
MCNs therefore construct what we call interpretive infrastructures—a system of informal yet durable knowledge that governs through belief, not verification/reality. Our findings contribute to CSCW’s growing attention to the epistemic and organizational dimensions of algorithmic labor. Folk theories, we argue, are not peripheral but are central to how algorithmic management operates in opaque, high-risk environments. They enable institutions to govern not through direct technical means, but through the circulation of strategically engineered narratives. This invites future research into how informal knowledge can become a durable, institutionally mediated mode of control in algorithmic workplaces.
6.3. Design and Policy Implications
Our findings suggest that designing for algorithmic systems in collaborative environments requires a shift in focus, from revealing how algorithms technically function to illuminating how algorithmic meaning is constructed, circulated, and institutionalized. In practice, workers interact not with raw code or backend logic, but with interpretations of the algorithm, often delivered and enforced through intermediary actors like MCNs or online communities  (Rosenblat and Stark, 2016; Bishop, 2020). Especially under conditions of opacity, these interpretations acquire operational power, shaping behavior and structuring labor.
Thus, design efforts should support what we call epistemic transparency: making visible not just the algorithm’s outcomes, but the infrastructures that produce and validate its meaning. Our study shows that such infrastructures are often embedded in mundane but powerful organizational practices. For instance, MCNs disseminate algorithmic “best practices” through scripted onboarding sessions and coaching routines, which not only instruct streamers on how to behave but also encode speculative beliefs about what the algorithm rewards. These beliefs are further materialized through equipment sales, where streamers are encouraged to purchase lighting, cameras, and décor upgrades as investments in algorithmic visibility. Rather than merely functional, these practices reflect and reinforce particular interpretations of the algorithm that benefit intermediary institutions while externalizing risk and responsibility onto workers. Designing for epistemic transparency thus requires surfacing such hidden infrastructures and the interests they serve, while also enabling alternative interpretations, such as folk theories emerging from peer networks, to support more pluralistic and contested understandings of algorithmic systems.
We also highlight epistemic asymmetry  (Ajmani et al., 2024, 2023) as a critical design and policy concern in algorithmic labor. Control over what is knowable about algorithms can be as powerful as control over the algorithm itself. To counter this, future systems might allow workers to annotate, challenge, or collaboratively construct algorithmic explanations, visualize competing interpretations, or track how dominant narratives shift over time. From a policy perspective, our findings call for stronger accountability mechanisms around algorithmic sensemaking. Platforms should not only disclose changes in algorithmic logic but also audit how such changes are communicated and operationalized by third parties such as MCNs. Regulatory frameworks might require platforms to monitor and report how algorithmic knowledge is intermediated, particularly when it becomes a basis for labor discipline or financial extraction. Policies could also mandate the provision of accessible explanatory resources, designed with input from workers and labor advocates, to mitigate dependency on opaque, potentially exploitative coaching ecosystems.
Finally, our work invites CSCW researchers, designers, and policymakers to view algorithmic sensemaking not as an individual cognitive task but as an organizational process. Institutions do not merely allocate labor or payment, and mediate access to meaning. Rather than only designing tools to increase algorithmic transparency, we invite CSCW and HCI scholars to ask: who gets to construct algorithmic meaning, for whom, and to what ends? Designing technologies for algorithmic collaboration therefore requires grappling with the political economy of epistemic authority and these questions raise broader design and policy challenges for shaping the future of algorithmic work.
7. Limitations and Future Work
As with all qualitative research, our findings are shaped by the specific contexts, participants, and interpretive lenses of our study. There are a few important limitations. First, this research draws on ethnographic fieldwork conducted in three MCNs in China, and while we prioritized diversity in location, streamer profile, and organizational structure, our sample cannot represent the full heterogeneity of the live-streaming industry, especially given regional, platform-specific, and organizational variation. Second, our focus on China brings unique sociopolitical and economic conditions that influence how platform labor and algorithmic governance unfold, including platform regulation, labor precarity, and the prominence of MCNs. While these conditions offer valuable insights into the global dynamics of platform intermediation, they may not be directly generalizable to other contexts. Third, our access to internal MCN practices was inevitably limited: despite prolonged immersion and trust-building, some dimensions of organizational decision-making and platform negotiation remained opaque or inaccessible. Finally, our analysis focuses on the interpretive and discursive construction of algorithmic knowledge, but does not directly evaluate the technical validity of the folk theories themselves or track their longitudinal impact on streamer outcomes – areas that future mixed-method or longitudinal studies could explore.
8. Conclusion
This study examined how algorithmic labor is governed not solely through platform design, but through the interpretive and organizational practices of intermediary institutions. Rather than focusing only on platform power or individual resistance, our analysis highlights how Multi-Channel Networks (MCNs) in the Chinese live-streaming industry play a central role in constructing and institutionalizing folk theories of algorithms. We identify a system of dual algorithmic theories: internally, MCNs acknowledge the volatility and opacity of platform systems, adopting probabilistic strategies to manage organizational risk; externally, they promote simplified, prescriptive narratives that portray the algorithm as transparent, fair, and responsive to individual effort. We show that these folk theories, though speculative, unverifiable and often inaccurate, can be mobilized as tools of labor control: providing streamers with simplified theories of algorithmic success, converting platform uncertainty into actionable scripts, and shifting responsibility for failure onto individuals. We call for future research to take seriously the role of intermediary institutions in algorithmic management and to attend to how folk theories of algorithms, when institutionalized, can shape the possibilities and limits of platform labor. By tracing how folk theories are used not only to explain uncertain algorithms, but to manage platform labor, this study reveals how interpretation itself becomes a key site of power in the algorithmic age.